OpenStudy (anonymous):

What is the uniform acceleration that causes a train’s velocity to change from 25 m/s to 75 m/s in an 5.0-s period? A. 20 m/s2 B. 5 m/s^2 C. 10 m/s2 D. 15 m/s^2

OpenStudy (anonymous):

@TerrelD99

OpenStudy (anonymous):

then the acceleration is 10 m/s2

OpenStudy (anonymous):

It doesn't tell me, but I got an 88% on my test so yeah..

OpenStudy (anonymous):

(75 - 25) / 5.0 = 50/5 = 10 meters per (.....)2

OpenStudy (anonymous):

and ok good !!

OpenStudy (anonymous):

Thanks ! (: